### v1

# Discovery of three magnetic He-sdOs with SALT

11institutetext: Institut für Physik und Astronomie, Universität Potsdam, Haus 28, Karl-Liebknecht-Str. 24/25, 14476 Potsdam-Golm, Germany 11email: dorsch@uni-potsdam.de 22institutetext: Armagh Observatory and Planetarium, College Hill, Armagh BT61 9DG, Northern Ireland 33institutetext: School of Mathematics and Physics, Queen’s University Belfast, Belfast, BT7 1NN, UK 44institutetext: Institute of Astronomy, The Observatories, Madingley Road, Cambridge CB3 OHA, UK 55institutetext: South African Astronomical Observatory, P.O. Box 9, Observatory Rd., Observatory 7935, Cape Town, South Africa 66institutetext: Department of Astronomy, University of Cape Town, Private Bag X3, Rondebosch 7701, South Africa 77institutetext: Australian Astronomical Optics, Faculty of Science and Engineering, Macquarie University, North Ryde, NSW 2113, Australia 88institutetext: Department of Physics, University of Nebraska at Omaha, 6001 Dodge St, Omaha, NE 68182-0266, USA
Discovery of three magnetic He-sdOs with SALT
M. Dorsch 11    C. S. Jeffery 22    A. Philip Monai 2233    C. A. Tout 44    E. J. Snowdon 2233    I. Monageng 5566    L. J. A. Scott 22    B. Miszalski 77    V. M. Woolf 88
(Received ; accepted )
Magnetic fields with strengths ranging from 300 to 500 kG have recently been discovered in a group of four extremely similar helium-enriched hot subdwarf (He-sdO) stars. Besides their strong magnetic fields, these He-sdO stars are characterised by common atmospheric parameters, clustering around $T_{\mathrm{eff}}$ = 46 500 K, $\log g/\mathrm{cm}\,\mathrm{s}^{-1}$ close to 6, and intermediate helium abundances. Here we present the discovery of three additional magnetic hot subdwarfs, J123359.44--674929.11, J125611.42--575333.45, and J144405.79--674400.93. These stars are again almost identical in terms of atmospheric parameters but, at $B\approx$ 200 kG, their magnetic fields are somewhat weaker than those previously known. The close similarity of all known He-sdOs implies a finely-tuned origin. We propose the merging of an He white dwarf with a H+He white dwarf. A differential rotation at the merge interface may initiate a toroidal magnetic field that evolves by a magnetic dynamo to produce a poloidal field. This field is either directly visible at the surface or may diffuse towards the surface if initially buried. We further discuss a broad absorption line centred at about 4630 Å that is common to all magnetic He-sdOs. This feature may not be related to the magnetic field but instead to the intermediate helium abundances in these He-sdO stars, allowing the strong He ii 4686 Å line to be perturbed by collisions with hydrogen atoms.
Key Words.:
stars: subdwarfs, stars: magnetic fields, line identification, —
1 Introduction
[…]
Here we present the discovery of three additional magnetic He-sdOs, J123359.44--674929.11, J125611.42--575333.45, and J144405.79--674400.93, hereinafter J1233--6749, J1256--5753, and J1444--6744, respectively [^2] J1233--6749 $\equiv$ Gaia DR3 5856712448201848320, J1256--5753 $\equiv$ Gaia DR3 6060271804721212672, J1444--6744 $\equiv$ Gaia DR3 5848132374839987840. . These stars closely mirror the atmospheric parameters of the known magnetic He-sdOs, but their average field strengths are slightly lower, between 180 and 220 kG. We also provide a list of six additional He-sdO stars that show a broad absorption feature at about 4630 Å – a feature that is also observed in all magnetic He-sdOs (Pelisoli et al., 2022). A potential non-magnetic origin of this spectral feature is briefly discussed.
2 Spectroscopic observations
[…]
To check for radial velocity and magnetic field variability, we obtained multiple SALT/RSS spectra of the new magnetic He-sdOs J1233--6749, J1256--5753 and J1444--6744. For each of these stars, the first (discovery) spectrum was taken in May 2023 and three follow-up spectra were taken at the end of January 2024, spaced by several days. A summary of these observations is given in Table 2.
Figure 1: He i 4471 Å and He ii 4542 Å for magnetic He-sdOs: previously known stars (top) and new ones from SALT/RSS (bottom). Observations are black and best-fit models red. All spectra are at rest wavelength, offset in steps of 0.3 in normalised flux. They were convolved to the RSS resolution (1.05 $\mathrm{\AA}$), except for that of J0415+2538, which has a worse resolution (1.9 $\mathrm{\AA}$).
Figure 2: Like Fig. 1, but for the unidentified feature, He ii 4686 Å and He i 4713 Å. Lorentz profile fits to the unidentified feature at about 4630 Å are shown in blue.
Table 1: Spectroscopic atmospheric parameters for confirmed magnetic He-sdOs (upper), new magnetic stars (centre), and He-sdOs showing the 4630 Å feature (lower). The last column lists the mean signal-to-noise ratio (S/N) of the RSS spectra. For J1233--6749, J1256--5753, J1444--6744, the four available spectra were co-added to obtain average magnetic field strengths while $T_{\mathrm{eff}}$, $\log g$, and $\log n(\mathrm{He})/n(\mathrm{H})$ were derived from the individual spectra. Estimated systematical uncertainties are stated for one sigma confidence.
Star $T_{\mathrm{eff}}$ / K $\log g/\mathrm{cm}\,\mathrm{s}^{-1}$ $\log n(\mathrm{He})/n(\mathrm{H})$ $B_{\mathrm{av}}$ / kG $v_{\mathrm{rad}}$ / km s-1 S/N
J0415+2528 $46580\pm 1500$ $5.98\pm 0.25$ $-0.10\pm 0.15$ $305\pm 20$ $-17\pm 10$
J0809--2627 $44900\pm 1000$ $5.93\pm 0.15$ $\phantom{+}0.28\pm 0.10$ $353\pm 10$ $\phantom{-}33\pm 2$
J1303+2646 $47950\pm 1500$ $5.97\pm 0.30$ $\phantom{+}0.25\pm 0.15$ $450\pm 20$ $-37\pm 8$
J1603+3412 $46450\pm 1500$ $6.06\pm 0.20$ $\phantom{+}0.07\pm 0.15$ $335\pm 15$ $\phantom{-0}6\pm 5$
J1233--6749 $47680\pm 1000$ $5.60\pm 0.13$ $\phantom{+}0.11\pm 0.05$ $222\pm 10$ $\phantom{-}82\pm 15$ 110
J1256--5753 $46980\pm 1000$ $5.75\pm 0.16$ $\phantom{+}0.00\pm 0.05$ $216\pm 24$ $\phantom{-0}5\pm 12$ 110
J1444--6744 $45990\pm 1000$ $5.55\pm 0.15$ $\phantom{+}0.17\pm 0.06$ $184\pm 14$ $\phantom{-0}1\pm 10$ 090
EC20577--5504 $46610\pm 1000$ $6.09\pm 0.12$ $\phantom{+}0.19\pm 0.05$ $\leq\phantom{0}59\pm\phantom{0}3$ $-42\pm 10$ 180
EC22332--6837 $46990\pm 1000$ $5.93\pm 0.12$ $\phantom{+}0.28\pm 0.06$ $\leq\phantom{0}82\pm\phantom{0}5$ $-101\pm 10$ 100
PG1625--034 $46570\pm 1100$ $5.67\pm 0.16$ $\phantom{+}0.49\pm 0.09$ $\leq 128\pm\phantom{0}9$ $-84\pm 11$ 040
J0714--2245 $44400\pm 1000$ $5.61\pm 0.13$ $-0.26\pm 0.05$ $\leq 107\pm\phantom{0}4$ $\phantom{+}33\pm 10$ 070
J1346--4025 $45070\pm 1000$ $5.83\pm 0.12$ $\phantom{+}0.13\pm 0.05$ $\leq\phantom{0}68\pm\phantom{0}4$ $-17\pm 10$ 130
J1946--4756 $46730\pm 1100$ $5.84\pm 0.16$ $\phantom{+}0.12\pm 0.08$ $\leq 129\pm\phantom{0}9$ $-41\pm 11$ 040
333 from Pelisoli et al. (2022). from Dorsch et al. (2022). J0714--2245 $\equiv$ Gaia DR3 2927944940172059648, J1346--4025 $\equiv$ UCAC4 248-062759, J1946--4756 $\equiv$ GALEX J194618.8-475617.
3 Spectral analysis
A direct way to identify a magnetic hot subdwarf is to resolve line splitting by the linear Zeeman effect. Hydrogen and helium lines are split into three components if the field is strong enough and appear broadened if the field is weaker. In addition, all of the previously known iHe-sdOH stars show a strong feature at about 4630 Å. The origin of this remains unidentified (Pelisoli et al., 2022, their fig. 5). We visually inspected the RSS spectra of all stars obtained as part of the SALT survey so far to identify possibly magnetic He-sdOs using both of these features.
We then used the model grid of Dorsch et al. (2022) to perform $\chi^{2}$ fits [^4] These fits were performed in the isis framework (Houck & Denicola, 2000) by the method of Irrgang et al. (2014). to these spectra and thus obtained best-fit magnetic field strengths, as well as atmospheric parameters $T_{\mathrm{eff}}$, $\log g$ and $\log n(\mathrm{He})/n(\mathrm{H})$. The best-fit parameters for all stars are listed in Table 1. These models are based on Tlusty (Hubeny & Lanz, 2017b) model atmospheres in non-local thermodynamic equilibrium. The inclusion of the magnetic field is simplistic: it is assumed to be homogeneous across the stellar surface and linear Zeeman splitting is only applied during the Synspec (Hubeny & Lanz, 2017a) spectral synthesis calculations [^5] The details are described in Appendix B of Dorsch et al. (2022).. While these models are not a realistic approximation of the actual field geometry, they are sufficient to detect Zeeman splitting and to estimate an average field strength.
3.1 Three new magnetic He-sdOs
Significant Zeeman splitting was detected in several He i lines and He ii 4686 Å for three stars, J1233--6749, J1256--5753 and J1444--6744, which we consider to be confirmed magnetic. Figs. 1 and 2 show the two most important spectral regions for these stars, He i 4471 Å and the range around He ii 4686 Å, respectively. The Zeeman splitting observed is weaker than in the four previously known magnetic He-sdOs reported by Dorsch et al. (2022) and Pelisoli et al. (2022), and is, in fact, close to the detection limit of RSS. However, there is no other effect that could explain the observed line shapes. The only conceivable other option would be rapid rotation at about 180 km s-1. As shown in Fig. 3, magnetic models provide a much superior fit to our RSS spectra when compared to non-magnetic rotating models. This is because rotation, unlike Zeeman splitting, can only increase the line broadening but not the line strengths.
Figure 3: Detailed view of He ii 4686 Å and He i 4713 Å for the new magnetic He-sdOs. The merged RSS spectra (black) are best matched by Zeeman-split models (red). Models at the best-fit atmospheric parameters but zero magnetic field (dashed red) are excluded even if $v\sin i$ is allowed to vary (blue). At zero field strength, the best-fit $v\sin i$ would equal 190, 180 and 130 km s-1 for J1233--6749, J1256--5753 and J1444--6744, respectively.
Like Pelisoli et al. (2022), we performed $\chi^{2}$ fits using two homogeneous magnetic field components that were allowed to vary in their field strength and contribution, expressed as a surface ratio. These $\chi^{2}$ fits resulted in the lowest average field strengths $B_{\mathrm{av}}$ in hot subdwarfs to date, $222\pm 10$ kG (J1233--6749), $216\pm 24$ kG (J1256--5753) and $184\pm 14$ kG (J1444--6744). These average field strengths were computed by taking the weighted mean of both components $B_{\mathrm{mean}}$ averaged over the four spectra available for each star, as stated in Table 2. Because our spectra were taken with one- and two-day gaps, variation in the magnetic field strength over these four spectra may result from a variable field geometry, possibly caused by rotation combined with magnetic spots. Individual fits are shown in Figs. 9 to 11 for the most important spectral lines; any variability in the field strengths should be confirmed with higher-quality spectra.
In contrast to the possibly variable fields, none of the three new magnetic He-sdOs showed signs of radial velocity variability, given the approximate 10 km s-1 accuracy achieved by RSS.
The magnetic fields found here are very likely non-homogeneous and more sophisticated models would be needed to further constrain their geometry. Such models would also benefit from a future determination of metal abundances. The strong N iii 4511, 4515 Å lines shown in Fig. 1 seem to suggest that the three new magnetic He-sdOs are significantly enriched in nitrogen compared to the Sun, especially J1444--6744. This star also seems to be enriched in neon as evidenced by strong Ne ii 4290, 4430 Å and Ne iii 4453, 4459 Å lines. Quantitative abundances will be measured by a more detailed analysis in the future.
Figure 4: Like Fig. 2, but for SALT/RSS spectra of stars that show the unidentified feature without significant Zeeman splitting.
3.2 The broad 4630 Å feature
[…]
Indeed, the origin of the 4630 Å feature [^6] It should be noted that the 4630 Å feature is in fact not always centred at a rest wavelength of exactly 4630 Å. For the iHe-sdOs considered here, its centre ranges between 4629 Å and 4635 Å at Lorentzian full width at half maxima between 20 Å and 38 Å. may be completely unrelated to the magnetic field. As already discussed by Pelisoli et al. (2022), this feature does not seem to be caused by Zeeman-split metal lines, diffuse interstellar bands, or even ultra-high excitation lines, a feature that is observed for some DO-type white dwarfs (Reindl et al., 2019). Instead, the feature 4630 Å feature seems to be related to the intermediate helium abundances: it could result from collisions with free hydrogen during the formation of He ii 4686 Å. This is similar to the blue satellite lines predicted for the $\mathrm{H}\alpha$ and $\mathrm{H}\beta$ lines in white dwarf stars by Allard et al. (2022) and Spiegelman et al. (2022). However, these features are not detected here, possibly because of the much higher $T_{\mathrm{eff}}$ and lower photospheric densities of intermediate He-sdO stars when compared to the white dwarfs considered by these authors.
4 Stellar parameters
Figure 5: Hertzsprung-Russell diagram showing known (dark blue) and new magnetic He-sdOs (light blue), as well as 4630Å-stars (violet). Merge tracks from Yu et al. (2021) for a metallicity of $Z=0.01$ and masses of 0.80, 0.65, 0.50 ${\rm M}_{\odot}$ are shown as black lines (solid: core helium burning, dashed: helium shell burning). The pre-helium main sequence phase is shown as a dotted line for 0.65 ${\rm M}_{\odot}$. The helium zero-age main sequence of Paczyński (1971) is shown in red and the grey region represents the EHB. Orange circles represent other He-sdOs from the SALT sample.
The spectral analysis performed in Sect. 3 provides atmospheric parameters, in particular the effective temperature and surface gravity. These can be combined with the Gaia DR3 (Gaia Collaboration et al., 2023) parallax $\varpi$ and an angular diameter $\Theta$ from photometric flux measurements to derive radii $R=\Theta/(2\varpi)$ and luminosities $L=4\pi R^{2}\sigma_{\mathrm{SB}}T^{4}_{\mathrm{eff}}$, where $\sigma_{\mathrm{SB}}$ is the Stefan–Boltzmann constant. To this end we performed spectral energy distribution (SED) fits to all He-sdOs discussed here. As usual, we used Gaia DR3 parallaxes with zero-point corrections applied according to Lindegren et al. (2021) and inflated uncertainties following El-Badry et al. (2021). We could also estimate stellar masses via $M=gR^{2}/G$. These remain poorly constrained due to the high uncertainty in the surface gravity $g$, resulting from additional line broadening introduced by non-homogeneous magnetic fields. To account for such systematic effects, we added estimated uncertainties of 1000 K in $T_{\mathrm{eff}}$, 0.15 in $\log g$ and 0.1 in $\log n(\mathrm{He})/n(\mathrm{H})$ in quadrature to the statistical uncertainties for the computation of stellar parameters. The resulting stellar parameters are listed in Table 3.
While the stellar masses are not well determined, the derived radii and luminosities are quite precise. All magnetic He-sdO stars are located close the zero-age helium main sequence (HeZAMS) in the Hertzsprung-Russell diagram (Fig. 5). This is not uncommon for He-sdO stars, as the comparison to the remaining SALT sample of He-sdO stars shows [^7] A detailed analysis of the full SALT sample will be published in a separate paper.. Indeed, the most likely formation scenario for all (non-binary) He-sdOs is the merging of two white dwarfs, including at least one helium-core WD (Han et al., 2002). The clustering of the now seven confirmed magnetic He-sdO stars at about 46 500 K on the HeZAMS further strengthens the proposition that these stars have formed through a very specific channel. If their masses are interpolated from their location on the HeZAMS of Paczyński (1971) in the Hertzsprung-Russell diagram, their mass distribution is centred at $M^{\mathrm{HeZAMS}}=0.85\pm 0.05$${\rm M}_{\odot}$, which is at the upper end of what is achievable by double HeWD merges (Zhang & Jeffery, 2012). This mass is a rough estimate given that (1) all magnetic He-sdOs retain significant amounts of hydrogen in their atmospheres while the HeZAMS assumes a pure helium composition and (2) the stars may also be somewhat evolved. The interpolated HeZAMS masses for all stars are also listed in Table 3.
5 Galactic kinematics
Galactic space velocities for the confirmed magnetic He-sdOs were calculated from their radial velocity (Table 1) and parallax and proper motion provided by Gaia DR3. Their Galactic orbits were computed with galpy [^8] http://github.com/jobovy/galpy, a python package for Galactic-dynamic calculations (Bovy, 2015). The resulting kinematic parameters are summarised in Table 4 while Fig. 12 shows the projection of the computed trajectories to the vertical distance ($z$) against Galactic radius ($R$) plane. The stars were roughly classified as Galactic thin disk, thick disk or halo stars by the method of Philip Monai et al. (2024). J1256--5753 and J1444--6744 are on almost circular orbits typical of thin disk stars while J1233--6749 is likely a thick disk star, given its larger eccentricity of $e=0.31\pm 0.01$. Similar calculations for the previously discovered magnetic He-sdOs show that, of the seven known magnetic stars, four are part of the thin disk, two are thick disk stars and one star is part of the Galactic halo, J1303+2646. This suggests that magnetic He-sdO stars can be formed in both young and old Galactic populations. This is consistent with the large range of possible delay times between the merging of double He-WD binaries and their initial formation as main sequence stars (Marsh et al., 1995; Yu & Jeffery, 2011).
6 Light curves
One might expect rotational light curve variability for strongly magnetic stars with inhomogeneous fields. Magnetic fields can cause dark or bright spots on a star’s surface, resulting in variations in brightness as the star rotates. For instance, bright magnetic spots are predicted to exist in O/B/A-type main sequence stars, caused by (sub-)surface convection zones (Cantiello & Braithwaite, 2011, 2019) or chemical spots (Shultz et al., 2018; Krticka et al., 2021), and are observed at low amplitudes in A- and late B-type stars (Balona et al., 2019). Similar variability was discovered for the helium-rich early B-type HD 144941 (Jeffery & Ramsay, 2018), a magnetic main sequence star (Przybilla et al., 2021; Shultz et al., 2021). Sub-surface convection zones are also predicted to exist in He-sdO stars (Groth et al., 1985), due to the ionisation of He ii. Momany et al. (2020) observed brightness variations in several hot subdwarfs in globular clusters at periods on the order of days and attributed this variability to rotational modulation caused by magnetic spots. Krtička et al. (2024) performed a spectroscopic analysis of three of these stars, ruling out binarity as the cause of the variations and detecting $T_{\mathrm{eff}}$ variability, though the presence of magnetic spots remains uncertain. Finally, flux variations in several cool magnetic white dwarfs are proposed to be caused by star spots (Brinkworth et al., 2013).
Considering the magnetic nature of the He-sdOs discussed here, the presence of magnetic spots on these stars seems at least plausible. We therefore searched for light curves in the Transiting Exoplanet Survey Satellite (TESS, Ricker et al., 2014) archive up to sector 70. Unfortunately, none of the new magnetic He-sdOs have been observed with TESS so far. However, TESS data are available for three of the stars showing the 4630 Å feature. We computed frequency-amplitude spectra for these stars.
EC20577--5504 was observed in TESS Sectors 27 and 67. It seems to show a highly significant variation at a period of 0.254 days and its first harmonic (Fig. 13: top). However, this variation is likely caused by the limited angular resolution of TESS images; the light curve seems to be contaminated by a close-by star (Gaia DR3 6458088855502103424).
J1346--4025 was observed in TESS Sectors 38 and 68 and shows an approximately sinusoidal modulation at a period of 5.4 days (Fig. 13: middle). This is commensurate with a rotational modulation and neither confirms nor excludes the presence of a magnetic field.
No light variation is detected for EC22332--6837 (TESS sectors 27, 28, 67, and 68). However, the S/N is such that the modulation seen in J1346--4025 would lie below the detection threshold for EC22332--6837 (Fig. 13: bottom).
7 Summary and discussion
We have identified three new magnetic He-sdO stars among the sample of 592 stars observed with SALT/RSS. This brings the total number of magnetic hot subdwarfs to seven stars that form a homogeneous class.
7.1 Fraction of magnetic stars
[…]
7.2 Magnetic field
All magnetic He-sdOs have observed field strengths of 200 to 500 kG and very similar surface properties with $T_{\mathrm{eff}}$ 46 500absent46500$\approx$ 46\,500≈ 46 500 K, $\log g\approx 6$ and intermediate helium abundances at $\log n(\mathrm{He})/n(\mathrm{H})$ $\approx 0$. That is to say, the magnetic sdO stars are very localised compared with the overall hot subdwarf population. A single-star origin seems unlikely for the following reasons.
  • 1)
    The subdwarf masses of about 0.8 ${\rm M}_{\odot}$ inferred from their locations on the helium main sequence are too high to have come from post-giant branch evolution, that is following common-envelope ejection or Roche-lobe stripping. This is because helium ignition occurs when a red-giant core reaches a mass of about 0.5 ${\rm M}_{\odot}$ and the envelope must be removed close to this point if the star is to become a hot subdwarf.
  • 2)
    Any pre-existing surface magnetic field would have been destroyed by the surface convection zone whilst the star was on the giant branch.
  • 3)
    Any pre-existing core magnetic field in a hot subdwarf would likely be destroyed by shell and core convection following helium ignition, as pointed out by Cantiello et al. (2016). While less convection would result from a non-degenerate He-ignition, this cannot lead to the formation of a single hot subdwarf star, given that there is no way to remove the hydrogen envelope.
  • 4)
    Differential rotation induced by contraction of the residual star from the giant branch to become a subdwarf could lead to a dynamo-driven field. This raises the question why this effect would not have occurred in all hot subdwarfs. There is no evidence for any detectable field in any other hot subdwarfs (Landstreet et al., 2012; Mathys et al., 2012; Randall et al., 2015).
A double-star origin seems possible. The most promising scenarios include a double He-WD merge or a merge of a hybrid HeCO-WD with a He-WD (Justham et al., 2011), as already discussed by Dorsch et al. (2022) and Pelisoli et al. (2022). Such merges are able to explain their helium-enriched surfaces and apparently single nature. This origin may be able to generate the magnetic fields but still needs to address the same question: why do not all massive subdwarfs have strong fields, in particular other He-sdO stars? A finely tuned formation scenario seems to be required.
Double-degenerate merge routes involving a fossil field in either progenitor were also considered.
  • 1)
    A fossil-field in the accretor fails because off-centre helium ignition creates convective shells after every shell flash, extending from close to surface, after the first shell flash, through to the core, once core burning is fully established. Flash-driven convection would progressively dissipate the field throughout the interior.
  • 2)
    A fossil-field in the donor fails because the donor material is fully mixed and/or fully convective at or near the merge. Hence a fossil field cannot survive a double white dwarf merge leading to the creation of a helium burning subdwarf.
This leaves the merging of two low-mass white dwarfs as the only realistic formation scenario for magnetic He-sdOs. Recently, Pakmor et al. (2024) conducted magnetohydrodynamical (MHD) simulations of the merging of two He-WDs with masses of either 0.3 + 0.25 ${\rm M}_{\odot}$ or 0.3 + 0.3 ${\rm M}_{\odot}$. They found that strong magnetic fields can be generated during both mergers. Their simulations indicate that if helium ignites off-center, strong convection will erase these fields. However, if helium ignites in the core, the fields can persist. This core ignition scenario is likely when the merging white dwarfs have similar core densities, which corresponds to nearly equal masses. While the observed magnetic He-sdOs appear to be more massive than those in the simulations, the underlying physics might be similar. Future research should use the results from these MHD simulations to develop post-merge evolutionary models, which can predict the properties of magnetic He-sdOs once they settle to the helium main sequence.
Because the helium enrichment in magnetic He-sdOs is less extreme than in the more common non-magnetic He-sdOs, new models that include a treatment of the magnetic field need to be able to retain a significant amount of surface hydrogen. Our proposal is that the magnetic subdwarfs arise from the merge of a He-WD with a H+He-WD. As they merge, the H+He-WD is destroyed and fully mixed. Most of its mass condenses onto the HeWD’s surface. Differential rotation (shear) exists at the interface between the accretor and the accreted material. This initiates a toroidal magnetic field. This toroidal field evolves by a magnetic dynamo to produce a poloidal field which is then visible at the surface.
A suitable double white dwarf progenitor would have a combined mass about 0.8 ${\rm M}_{\odot}$, with component masses of about 0.4 to 0.5 ${\rm M}_{\odot}$ for the He-WD and about 0.3 to 0.4 ${\rm M}_{\odot}$ for the H+He-WD. The H+He-WD could be marginally the more massive of the pair, so long as it is has the larger radius. The hydrogen content of either component cannot be too large because the location of iHe-sdOH stars on the helium main sequence requires that there be negligible hydrogen beneath the atmosphere. Potential systems are identified theoretically (Han, 1998) and observed in both the Galactic disk and halo (Brown et al., 2020). They are believed to arise as a consequence of two common-envelope ejection episodes in an intermediate mass binary, where both episodes occur when each donor star is on the red-giant branch. The condition that some hydrogen remains on the surface of either He-WD is that the common-envelope component shrinks more quickly than the orbit decays, and before all of the hydrogen is completely removed. Hall & Jeffery (2016) discussed similar cases for a combined mass of about 0.5 ${\rm M}_{\odot}$ and found very little hydrogen could survive. An important feature of the merging double white dwarf model is that the first helium-shell flash drives a strong convection zone upwards. If this reaches the surface it reveals carbon and removes residual hydrogen (Zhang & Jeffery, 2012). Such a convective pulse would also redistribute angular momentum from the original shear layer, moving the shear layer towards the surface. The absence of carbon in the spectra of magnetic He-sdOs implies that complete mixing has not occurred and that a field-generating shear layer could still exist beneath the stellar surface. Such constraints may contribute in part to the fine tuning required to model the iHe-sdOH stars.
This proposal needs to be tested more fully in a number of ways.
  • 1)
    Demonstrate that the merge leads to sufficient differential rotation in the mixed envelope to give rise to a strong magnetic field.
  • 2)
    Determine properties, such as mass, residual hydrogen, and differential rotation, necessary for the merged system to generate a dynamo by population synthesis or other means.
  • 3)
    Establish the progenitor properties at the merge necessary to yield 2), such as He-WD mass, H+He-WD mass and H+He-WD core mass.
It is likely that all of these issues need to be resolved if we are to understand why there is a small sweet spot for generating strong magnetic fields in hot subdwarfs. The proposed merging of a HeWD with H+He-WD by itself can not explain why the observed magnetic He-sdOs are so similar. However, Pakmor et al. (2024) suggest that only white dwarfs with similar core densities can maintain a strong surface magnetic field in the He-sdO remnant. This finding may explain the observed clustering in the HRD because similar core densities in the merging WDs would result in similar core masses for the He-sdOs.
Unfortunately, the detailed surface metal abundances of the magnetic He-sdOs discovered here are not known because of the limited quality of the available spectra. Far-UV spectra taken with the Hubble Space Telescope could solve this issue but are presently not available.
7.3 Kinematics and light curves
We further performed Galactic orbit calculations for all seven known magnetic He-sdO stars. Given that we find four stars in the thin disk, two in the thick disk and one in the Galactic halo, magnetic He-sdOs seem to exist in all Galactic populations, both old and young. Unfortunately, no TESS light curves are available for the three new He-sdOHs. Light curve variability therefore remains undetected for all seven He-sdOHs and follow-up time series photometry should be performed.
7.4 4630Å absorption
[…]
Acknowledgements.
We thank N. F. Allard for her useful comments on the 4630 Å feature and the anonymous referee for their detailed comments. MD is supported by the Deutsches Zentrum für Luft- und Raumfahrt (DLR) through grant 50-OR-2304. CAT thanks Churchill College for his fellowship. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This paper includes data collected by the TESS mission, which are publicly available from the Mikulski Archive for Space Telescopes (MAST). Funding for the TESS mission is provided by NASA’s Science Mission directorate. This research has made use of NASA’s Astrophysics Data System.
Appendix A Additional material
Table 2: Summary table of the RSS spectra taken of the three new magnetic He-sdOs. The last column lists the weighted average field strength $B_{\mathrm{mean}}$ of the two-component best-fit for each spectrum, as discussed in Sect. 3.1. $A_{2}/A_{1}$ is the surface ratio of the second field component ($B_{2}$) relative to the first ($B_{1}$).
Star JD $t_{\mathrm{exp}}/\mathrm{s}$ S/N $B_{1}/\mathrm{kG}$ $B_{2}/\mathrm{kG}$ $A_{2}/A_{1}$ $B_{\mathrm{mean}}/\mathrm{kG}$ ID
J1233--6749 2460069.27553 600 81 $178\pm 4$ $625\pm 26$ $0.14\pm 0.02$ $232\pm\phantom{0}6$ a
2460337.55751 600 43 $178\pm 8$ $550\pm 50$ $0.14\pm 0.04$ $225\pm 11$ b
2460340.59014 600 95 $178\pm 4$ $568\pm 30$ $0.15\pm 0.02$ $227\pm\phantom{0}7$ c
2460341.50690 600 35 $165^{+70}_{-30}$ $250^{+50}_{-70}$ $0.87^{+13}_{-0.7}$ $204\pm 42$ d
J1256--5753 2460069.27480 500 31 $260^{+25}_{-170}$ $130^{+30}_{-40}$ $0.77^{+0.61}_{-0.46}$ $201\pm 44$ a
2460334.54060 500 67 $164\pm 8$ $318\pm 16$ $0.43\pm 0.10$ $211\pm 12$ b
2460336.52242 500 78 $140\pm 8$ $270\pm 10$ $0.76\pm 0.13$ $196\pm\phantom{0}8$ c
2460341.49358 500 60 $212\pm 4$ $625\pm 50$ $0.12\pm 0.02$ $258\pm\phantom{0}8$ d
J1444--6744 2460071.40057 500 48 $147\pm 10$ $277\pm 13$ $0.73\pm 0.16$ $202\pm\phantom{0}9$ a
2460364.52464 500 35 $143\pm 13$ $310\pm 37$ $0.25\pm 0.09$ $189\pm 14$ b
2460369.51603 500 45 $133\pm 15$ $248\pm 45$ $0.38\pm 0.29$ $165\pm 25$ c
2460371.57299 500 67 $145\pm 6$ $295\pm 16$ $0.31\pm 0.05$ $180\pm\phantom{0}7$ d
Table 3: Best stellar parameters based on the combination of spectroscopic atmospheric parameters from Table 1, the Gaia parallax, and the SED’s angular diameter. Estimated systematic uncertainties of 1000 K in $T_{\mathrm{eff}}$, 0.15 in $\log g$, and 0.1 in $\log n(\mathrm{He})/n(\mathrm{H})$ were added in quadrature to the statistical uncertainties for the computation of the radii $R$, luminosities $L$, and masses $M$. Stellar parameters are stated as median values with one sigma uncertainties. The last column states masses obtained by interpolation in the helium main sequence of Paczyński (1971), given the uncertainties on $T_{\mathrm{eff}}$ and $L$.
Star $\log\Theta/\mathrm{rad}$ $E(44$--$55)/\mathrm{mag}$ $\varpi/\mathrm{mas}$ $R$ / ${\rm R}_{\odot}$ $L$ / ${\rm L}_{\odot}$ $M^{\mathrm{spec}}$ / ${\rm M}_{\odot}$ $M^{\mathrm{HeZAMS}}$ / ${\rm M}_{\odot}$
J0415+2528 $-11.476\pm 0.006$ $0.298\pm 0.004$ $0.49\pm 0.06$ $0.151^{+0.020}_{-0.016}$ $\phantom{0}97^{+31}_{-22}$ $0.80^{+0.60}_{-0.40}$ $0.82\pm 0.06$
J0809--2627 $-11.246\pm 0.005$ $0.072\pm 0.003$ $0.68\pm 0.04$ $0.185^{+0.011}_{-0.010}$ $125^{+19}_{-17}$ $1.06^{+0.31}_{-0.24}$ $0.82\pm 0.03$
J1303+2646 $-11.517\pm 0.006$ $0.005\pm 0.003$ $0.33\pm 0.07$ $0.210^{+0.050}_{-0.040}$ $200^{+120}_{-70}$ $1.40^{+2.40}_{-0.70}$ $0.90\pm 0.07$
J1603+3412 $-11.806\pm 0.006$ $0.027\pm 0.005$ $0.22\pm 0.07$ $0.160^{+0.070}_{-0.040}$ $110^{+120}_{-50}$ $1.00^{+2.20}_{-0.50}$ $0.83\pm 0.06$
J1233--6749 $-11.270\pm 0.004$ $0.21\pm 0.01$ $0.66\pm 0.04$ $0.181^{+0.009}_{-0.009}$ $153^{+21}_{-19}$ $0.48^{+0.22}_{-0.16}$ $0.88\pm 0.04$
J1256--5753 $-11.190\pm 0.005$ $0.25\pm 0.01$ $0.82\pm 0.03$ $0.175^{+0.007}_{-0.007}$ $129^{+16}_{-14}$ $0.74^{+0.34}_{-0.24}$ $0.85\pm 0.03$
J1444--6744 $-11.141\pm 0.005$ $0.36\pm 0.01$ $0.77\pm 0.05$ $0.210^{+0.015}_{-0.013}$ $173^{+31}_{-25}$ $0.56^{+0.27}_{-0.19}$ $0.86\pm 0.04$
EC20577--5504 $-11.542\pm 0.004$ $0.03\pm 0.01$ $0.54\pm 0.07$ $0.117^{+0.017}_{-0.013}$ $\phantom{0}58^{+19}_{-13}$ $0.62^{+0.36}_{-0.23}$ $0.80\pm 0.04$
EC22332--6837 $-11.584\pm 0.005$ $0.04\pm 0.01$ $0.33\pm 0.05$ $0.177^{+0.028}_{-0.022}$ $140^{+50}_{-40}$ $1.00^{+0.60}_{-0.40}$ $0.86\pm 0.04$
PG1625--034 $-11.347\pm 0.003$ $0.23\pm 0.01$ $0.59\pm 0.05$ $0.168^{+0.014}_{-0.012}$ $122^{+25}_{-20}$ $0.54^{+0.27}_{-0.18}$ $0.84\pm 0.04$
J0714--2245 $-11.180\pm 0.005$ $0.15\pm 0.01$ $0.76\pm 0.03$ $0.192^{+0.008}_{-0.008}$ $130^{+17}_{-15}$ $0.48^{+0.24}_{-0.16}$ $0.82\pm 0.03$
J1346--4025 $-11.000\pm 0.005$ $0.07\pm 0.01$ $1.43\pm 0.05$ $0.156^{+0.006}_{-0.006}$ $\phantom{0}91^{+11}_{-10}$ $0.58^{+0.27}_{-0.18}$
J1946--4756 $-11.690\pm 0.005$ $0.04\pm 0.01$ $0.29\pm 0.08$ $0.160^{+0.060}_{-0.040}$ $110^{+90}_{-40}$ $0.61^{+0.60}_{-0.28}$ $0.84\pm 0.05$
999 Originally from Pelisoli et al. (2022). Originally from Dorsch et al. (2022). Too far away from the helium main sequence for a sensible mass estimate.
Figure 6: The co-added RSS spectrum of J1233--6749 (black) compared to the best two-component model (red).
Figure 7: The co-added RSS spectrum of J1256--5753 (black) compared to the best two-component model (red).
Figure 8: The co-added RSS spectrum of J1444--6744 (black) compared to the best two-component model (red).
Figure 9: Best-fits (red) to individual spectra (black) of J1233--6749. Each spectrum is identified by a letter as listed in Table 2. Only the most important lines are shown. The models were allowed to vary in magnetic field strength, component contribution, and radial velocity. Other atmospheric parameters were kept the same for all spectra.
Figure 10: Like Fig. 9, but for J1256--5753.
Figure 11: Like Fig. 9, but for J1444--6744.
Figure 12: Our galpy orbits for the confirmed magnetic He-sdOs, showing motion in the $R$ versus $z$ plane integrated over 3 Gyrs from their current position. The top row shows the stars discussed in this paper while the bottom row shows the previously discovered magnetic He-sdO stars.
Table 4: Galactic velocities of the new and the known magnetic He-sdOs and orbital parameters including the $z$-component of angular momentum ($L_{z}$), eccentricity (e), galactic apocentre ($R_{\rm apo}$), galactic pericentre ($R_{\rm peri}$) and maximum vertical amplitude ($z_{\rm max}$). The left-handed system for the Galactic velocity components is used here, where $U$ is positive towards the Galactic Centre, $V$ is in the direction of Galactic rotation and $W$ is positive towards the North Galactic Pole. The last column shows the population classification as described in Philip Monai et al. (2024), where TH $\equiv$ thin disk, TK $\equiv$ thick disk, and H $\equiv$ halo.
Star $U$ /  $\mbox{km}\,\mbox{s}^{-1}$ $V$ /  $\mbox{km}\,\mbox{s}^{-1}$ $W$ /  $\mbox{km}\,\mbox{s}^{-1}$ $L_{z}$ / kpc  $\mbox{km}\,\mbox{s}^{-1}$ $e$ $R_{\mathrm{apo}}$ / kpc $R_{\mathrm{peri}}$ / kpc $z_{\rm max}$ / kpc pop
J1233-6749 $-84.5$$\pm$ $3.4$ $175.0$$\pm$ $8.6$ $-51.0$$\pm$ $1.0$ $1374$$\pm$ $65$ $0.36$$\pm$ $0.03$ $8.72$$\pm$ $0.01$ $4.12$$\pm$ $0.30$ $1.02$$\pm$ $0.04$ TK
J1256-5753 $-52.1$$\pm$ $3.0$ $242.0$$\pm$ $6.0$ $-4.0$$\pm$ $0.6$ $1858$$\pm$ $47$ $0.11$$\pm$ $0.01$ $8.51$$\pm$ $0.11$ $6.90$$\pm$ $0.28$ $0.18$$\pm$ $0.01$ TH
J1444-6744 $-49.3$$\pm$ $3.3$ $226.2$$\pm$ $4.5$ $16.2$$\pm$ $0.7$ $1536$$\pm$ $38$ $0.19$$\pm$ $0.02$ $7.66$$\pm$ $0.04$ $5.20$$\pm$ $0.21$ $0.15$$\pm$ $0.01$ TH
J0415+2538 $-18.7$$\pm$ $5.4$ $256.6$$\pm$ $0.9$ $16.6$$\pm$ $1.8$ $2695$$\pm$ $40$ $0.13$$\pm$ $0.01$ $13.44$$\pm$ $0.30$ $10.23$$\pm$ $0.17$ $0.88$$\pm$ $0.07$ TH
J0809--2627 $-29.5$$\pm$ $0.8$ $231.2$$\pm$ $1.0$ $11.2$$\pm$ $0.1$ $2099$$\pm$ $10$ $0.07$$\pm$ $0.01$ $9.37$$\pm$ $0.05$ $8.13$$\pm$ $0.05$ $0.16$$\pm$ $0.00$ TH
J1303+2646 $-110.1$$\pm$ $5.5$ $363.0$$\pm$ $6.2$ $49.7$$\pm$ $6.1$ $1729$$\pm$ $50$ $0.53$$\pm$ $0.05$ $17.50$$\pm$ $1.93$ $5.39$$\pm$ $0.18$ $7.79$$\pm$ $1.83$ H
J1603+3412 $-25.7$$\pm$ $1.2$ $255.4$$\pm$ $2.6$ $101.0$$\pm$ $7.3$ $1173$$\pm$ $130$ $0.35$$\pm$ $0.07$ $9.01$$\pm$ $0.32$ $4.36$$\pm$ $0.51$ $4.31$$\pm$ $0.68$ TK
Figure 13: TESS light curves (left) and frequency-amplitude spectra (right) for three He-sdOs that show the 4630 Å feature. Light curves in units of relative brightness are shown for the best single sector in each case (as labelled). Times are in Julian days after start of the TESS mission. The amplitude spectra in units of relative brightness were computed from all available sectors. The inset shows the spectral window function at the same horizontal scale. Note that the light curve of EC20577--5504 is likely contaminated by a close-by star.